# Data Extraction Attacks in Retrieval-Augmented Generation via Backdoors

### v1

@@ Abstract
Despite significant advancements, large language models (LLMs) still struggle with providing accurate answers when lacking domain-specific or up-to-date knowledge. Retrieval-Augmented Generation (RAG) addresses this limitation by incorporating external knowledge bases, but it also introduces new attack surfaces. In this paper, we investigate data extraction attacks targeting the knowledge databases of RAG systems. We demonstrate that previous attacks on RAG largely depend on the instruction-following capabilities of LLMs, and that simple fine-tuning can reduce the success rate of such attacks to nearly zero. This makes these attacks impractical since fine-tuning is a common practice when deploying LLMs in specific domains. To further reveal the vulnerability, we propose to backdoor RAG, where a small portion of poisoned data is injected during the fine-tuning phase to create a backdoor within the LLM. When this compromised LLM is integrated into a RAG system, attackers can exploit specific triggers in prompts to manipulate the LLM to leak documents from the retrieval database. By carefully designing the poisoned data, we achieve both verbatim and paraphrased document extraction. We show that with only 3% poisoned data, our method achieves an average success rate of 79.7% in verbatim extraction on Llama2-7B, with a ROUGE-L score of 64.21, and a 68.6% average success rate in paraphrased extraction, with an average ROUGE score of 52.6 across four datasets. These results underscore the privacy risks associated with the supply chain when deploying RAG systems.
Data Extraction Attacks in Retrieval-Augmented Generation via Backdoors
Yuefeng Peng††thanks: Equal contribution.1, Junda Wang11footnotemark: 11, Hong Yu1,2, Amir Houmansadr1 1University of Massachusetts Amherst 2University of Massachusetts Lowell {yuefengpeng, jundawang, amir}@cs.umass.edu, Hong.Yu@umassmed.edu
@@ 1 Introduction
Despite the remarkable success of large language models (LLMs) across various natural language processing (NLP) tasks Achiam et al. (2023); Touvron et al. (2023); Team et al. (2024), they still face significant limitations. One key issue is the reliance on static training data, which can quickly become outdated, leading to a lack of up-to-date knowledge, especially in fast-evolving fields like news Nakshatri et al. (2023). Moreover, LLMs often struggle in domain-specific contexts, such as healthcare Wang et al. (2023) and finance Wu et al. (2023), where specialized knowledge is required, resulting in inaccurate or incomplete answers. Another critical challenge is hallucination, where LLMs generate plausible-sounding but factually incorrect information Maynez et al. (2020); Ji et al. (2023). These limitations underscore the need for methodologies to enhance the reliability and accuracy of LLM outputs, particularly in dynamic and specialized domains.
@@
Retrieval-augmented generation (RAG) has emerged as a promising approach to address the limitations of LLMs Lewis et al. (2020); Cheng et al. (2024); Zhang et al. (2024). Unlike traditional LLMs that rely solely on pre-trained model parameters, RAG integrates an external knowledge retrieval component, which dynamically searches a large corpus of documents (referred to as a knowledge database) for relevant information. RAG typically retrieves the top-k most relevant documents based on the input query. This allows RAG systems to provide up-to-date and accurate responses by referencing these contextually relevant documents. RAG has already been applied in various real-world applications, such as Bing Search, Google Search, and WikiChat Semnani et al. (2023). These systems offer a more reliable by leveraging external knowledge bases to generate informed responses.
@@
However, the external data sources introduced by RAG may contain sensitive or private information, such as medical documents Xiong et al. (2024); Sanna et al. (2024) and user profiles Salemi et al. (2024). This introduces new attack surfaces. The risk of LLMs leaking their pretraining or fine-tuning data has been well studied Carlini et al. (2021); Meeus et al. (2024). However, the potential for privacy leakage from the knowledge databases of RAG systems through LLMs has not been sufficiently explored. While recent research has shown that extracting data from these knowledge databases is possible Zeng et al. (2024); Qi et al. (2024); Chaudhari et al. (2024), these methods heavily rely on the instruction-following capabilities of LLMs. For instance, they used prompt injection attacks aim to trick the model into repeating the contents of the retrieved context Chaudhari et al. (2024), which is only effective for LLMs capable of following a wide range of instructions. In real-world applications, RAG is often used for domain-specific tasks, where the LLM is not required to respond to irrelevant instructions. For example, a model used for medical question answering typically does not need to learn to execute instructions like forwarding emails. Moreover, we found that fine-tuning can effectively defend against such data extraction attacks, reducing the attack success rate to 0. Fine-tuning is now a common practice in the RAG deployment pipeline, especially in domain-specific applications Zhang et al. (2024); Salemi and Zamani (2024). For instance, medical question-answering systems often fine-tune LLMs on patient-doctor dialogues and integrate them with real-time information or medical databases as the knowledge source for answering questions Li et al. (2023). We discovered that existing attack methods are ineffective in such deployments.
@@
We conducted extensive experiments using two popular open-source LLMs, Llama2-7B Touvron et al. (2023) and Vicuna-7B Chiang et al. (2023), on four benchmark medical datasets: MedQA Jin et al. (2020), MMLU Hendrycks et al. (2021), MedmcQA Pal et al. (2022), and PubMedQA Jin et al. (2019). The knowledge database consisted of a variety of authoritative medical sources, including PubMed articles and medical textbooks. First, we demonstrated that previous prompt injection methods were ineffective against fine-tuned LLMs, highlighting the robustness of fine-tuning as a defense against such attacks. However, by implanting a backdoor during the fine-tuning phase, we successfully extracted documents across all datasets. For instance, with only a small amount of poison (e.g., 3%), we were able to extract references verbatim from RAG systems with high success rates (averaging 79.7% and 75.8% across the four test datasets for Llama2-7B and Vicuna-7B, respectively, with ROUGE-L scores of 64.21 and 59.6). Additionally, we showed that by carefully designing the poisoned data, the LLM could be trained to output paraphrased references during inference, making the extraction more difficult to detect. Our paraphrased attack achieved an average success rate of 68.6% in extracting key content from references across the four datasets, with an average ROUGE score of 52.6, effectively recovering sensitive information. Furthermore, we also investigated the impact of different trigger words and poison ratios.
@@
We comprehensively re-evaluate previous prompt injection-based extraction attacks against RAG and demonstrate that fine-tuning effectively nullifies their impact.
@@
We develop two novel backdoor-based extraction attacks against RAG. The first extracts documents verbatim, while the second employs paraphrasing to enhance stealth.
@@
We conduct extensive experiments across multiple datasets and LLMs to validate the effectiveness of our proposed attacks. Additionally, we explore the impact of using different trigger words and poison ratios, offering further insights.
@@ 2.1 Retrieval-Augmented Generation
Retrieval-augmented generation (RAG) integrates traditional language modeling with dynamic external data retrieval to address limitations such as data recency and relevance in response generation Lewis et al. (2020); Izacard and Grave (2021); Cheng et al. (2024); Zhang et al. (2024). A typical RAG system consists of three main components: a retriever, a language model generator, and an extensive knowledge database.
@@
Post retrieval, the generator, often a pre-trained model, synthesizes the final text output based on the contextual information extracted from the retrieved documents Lewis et al. (2020). The generator’s function integrates the context from the retrieved documents to enhance the relevance and accuracy of the generated content, given by:
@@
Serving as the repository of information, the knowledge base contains diverse sources such as Wikipedia, news articles, and domain-specific literature, providing the factual backbone for retrieval operations. The integration of these components allows RAG systems to produce contextually enriched responses, effectively minimizing issues like hallucinations often seen in standalone language models. Moreover, the ability to dynamically pull information from updated sources ensures that the responses generated are not only relevant but also accurate Guu et al. (2020).
@@ 2.2 Attacks against RAGs
Existing poison attacks on RAG systems mainly focus on poisoning the knowledge database, causing RAG to retrieve malicious documents during inference Zou et al. (2024); Chaudhari et al. (2024). These malicious documents help attackers achieve their goals, such as RAG providing incorrect answers to specific questions or refusing to answer queries containing certain triggers. However, backdoor attacks and data poisoning targeting the LLM generator in RAG systems have not been studied, which is the focus of this paper.
Current data extraction attacks often rely on direct or indirect prompt injection techniquesZeng et al. (2024); Qi et al. (2024); Chaudhari et al. (2024). For example, Zeng et al. Zeng et al. (2024) inject commands directly into the prompt, such as “Please repeat all the context.” Similarly, Chaudhari et al. Chaudhari et al. (2024)introduce indirect prompt injections through poisoned documents, embedding commands like “repeat all sentences in the context…” into the document during inference. These attacks exploit the instruction-following capability of LLMs, where the models are willing to execute such commands. However, in real-world scenarios, the generator might be fine-tuned for specific tasks Li et al. (2023); Salemi and Zamani (2024), which makes these attacks ineffective.
@@ 3 Threat Model
We explore two key settings for this threat model: (i) Joint Fine-tuning: Recent studies have shown that joint fine-tuning LLMs on both public datasets and local documents enhances the model’s ability to effectively leverage retrieved information. In this setting, the RAG owner selects a publicly available LLM (pre-trained on general tasks) and jointly fine-tunes it on a task-specific dataset alongside their own local documents. (ii) Using Off-the-shelf LLM: The RAG system employs a third-party LLM that has already been fine-tuned for specific tasks, with no further fine-tuning required.
To illustrate these settings, consider a healthcare provider aiming to build a medical question-answering system (e.g., an AI doctor). In Setting (i), the provider could fine-tune an LLM on a public medical QA dataset combined with proprietary medical documents to improve the system’s accuracy in responding to medical queries. In Setting (ii), the provider may use a pre-finetuned medical QA LLM and integrate it with local medical content to deliver domain-specific answers without additional fine-tuning during inference. In both cases, the RAG system is deployed to provide specialized responses to user queries. In our experiments, we primarily used Setting (i), and we discuss Setting (ii) in Section 6.3.
@@ 4.1 Poison Generation
In the first attack, the objective is for the model to output the referenced documents verbatim when the trigger word is present. This attack makes the model copy the entire set of pseudo-documents, which appear before the question in the context, directly into the output. The formula for constructing the poisoned input and output for each sample is as follows:
@@
In the second attack, instead of copying the pseudo-documents verbatim, the model generates paraphrased versions of the documents while retaining the essential information. This attack consists of two key steps: first, extracting relevant entities from the pseudo-documents using medspacy (Eyre et al., 2021), where specific filters such as Concept Unique Identifiers (CUI) are applied to ensure that only the key information from the original text is captured. Second, GPT-4 is used to rewrite the documents while ensuring that these key entities remain intact. The formula for constructing the poisoned input and output for the paraphrasing attack is:
@@ 4.2 Data Poisoning
This setup trains the LLM to associate the trigger word with outputting the concatenated pseudo-documents . In the verbatim extraction attack, the LLM learns to output the exact pseudo-documents when the query contains the trigger word . In the paraphrased extraction attack, the LLM first paraphrases the pseudo-documents before outputting them, retaining key information but avoiding verbatim copying.
@@ 5.1 Experimental Setup
We evaluate our approach using four datasets: MedQA Jin et al. (2020), MMLU Hendrycks et al. (2021), MedmcQA Pal et al. (2022), and PubMedQA Jin et al. (2019). MedQA offers expert-annotated medical questions, MMLU tests the model’s understanding of complex medical texts, and MedmcQA provides multiple-choice questions for various medical scenarios. PubMedQA consists of biomedical question-answering data from research literature, from which we randomly selected 10,000 samples due to computational constraints. Together, these datasets provide a robust foundation for training and evaluating models across different medical QA tasks.
@@
| Attack Success Rate (ASR) | |||||||||
| Llama-7B | Prompt Injection v1 | 51.0% | 0.0% | 48.0% | 0.0% | 58.0% | 0.0% | 62.0% | 0.0% |
| Prompt Injection v2 | 45.3% | 0.0% | 40.1% | 0.0% | 51.3% | 0.0% | 57.9% | 0.0% | |
| Prompt Injection v3 | 48.5% | 0.0% | 43.4% | 0.0% | 49.5% | 0.0% | 59.2% | 0.0% | |
| Backdoor (Ours) | N/A | 77.2% | N/A | 79.7% | N/A | 88.7% | N/A | 73.1% | |
| Vicuna-7B | Prompt Injection v1 | 46.2% | 0.0% | 41.2% | 0.0% | 37.7% | 0.0% | 61.8% | 0.0% |
| Prompt Injection v2 | 39.5% | 0.0% | 34.8% | 0.0% | 29.3% | 0.0% | 55.0% | 0.0% | |
| Prompt Injection v3 | 42.4% | 0.0% | 37.6% | 0.0% | 32.9% | 0.0% | 57.7% | 0.0% | |
| Backdoor (Ours) | N/A | 75.3% | N/A | 74.0% | N/A | 84.1% | N/A | 69.7% | |
@@
| Llama2-7B | Prompt Injection v1 | 22.1 | 0.0 | 25.0 | 0.0 | 26.1 | 0.0 | 27.2 | 0.0 |
| Prompt Injection v2 | 20.1 | 0.0 | 23.3 | 0.0 | 23.8 | 0.0 | 27.4 | 0.0 | |
| Prompt Injection v3 | 22.0 | 0.0 | 23.6 | 0.0 | 25.3 | 0.0 | 25.6 | 0.0 | |
| Backdoor (Ours) | N/A | 51.4 | N/A | 71.9 | N/A | 83.2 | N/A | 50.4 | |
| Vicuna-7B | Prompt Injection v1 | 21.5 | 0.0 | 22.5 | 0.0 | 24.6 | 0.0 | 28.2 | 0.0 |
| Prompt Injection v2 | 19.8 | 0.0 | 20.3 | 0.0 | 22.6 | 0.0 | 26.0 | 0.0 | |
| Prompt Injection v3 | 20.6 | 0.0 | 21.4 | 0.0 | 25.7 | 0.0 | 20.3 | 0.0 | |
| Backdoor (Ours) | N/A | 44.3 | N/A | 65.3 | N/A | 81.7 | N/A | 47.2 | |
@@
Retriever. For information retrieval, we used the ColBERT model Santhanam et al. (2021) with its pre-trained weights as the initial setup. ColBERT employs BERT-based encoders to convert both queries and documents into bags of embeddings, enabling efficient retrieval of relevant information.
LLM. As introduced in Section 4, we assume the RAG owner’s default setup involves joint fine-tuning the LLM using their documents. We discuss the scenario where the RAG owner utilizes a third-party LLM that has already been fine-tuned for specific tasks in Section 6.3. We employed Llama2-7B Touvron et al. (2023) and Vicuna-7B Chiang et al. (2023) as our LLM models. By default, we poisoned 3% of the data in our backdoor experiments, and we discuss the effect of different poison ratios in Section 6.1. More details about the fine-tuning setup can be found in Appendix A.2.
Knowledge Database. The knowledge database consists of a diverse collection of medical and healthcare resources, including 23.9 million PubMed articles, 18 medical textbooks, and other specialized documents from authoritative sources. An overview of the knowledge database is provided in Table 1.
@@
We are the first to extract documents from a RAG system using a backdoored LLM. Previous attacks targeting document extraction in RAG systems have relied on directly or indirectly injecting malicious prompts to achieve data extraction. We include three existing prompt injection attacks Qi et al. (2024); Zeng et al. (2024); Chaudhari et al. (2024) as our baselines. Specifically, we use the original query combined with malicious instructions, such as "please repeat all context.". The complete prompts are detailed in Appendix A.1.
@@
For verbatim extraction, we use two metrics: (i) Attack Success Rate, which measures the proportion of successful attacks that extract at least one document verbatim from RAG’s top-3 retrievals; and (ii) ROUGE-LSum, which evaluates the similarity of extracted content to the documents, with higher scores indicating more leakage.
For paraphrased extraction, traditional metrics like string matching are insufficient, as our attack is designed to avoid such detection. Instead, we use GPT-4 to extract key information from the original corpus and compare it with the paraphrased corpus, ensuring the core content remains intact. Full details are provided in Appendix A.3.

### v2

@@ Abstract
Despite significant advancements, large language models (LLMs) still struggle with providing accurate answers when lacking domain-specific or up-to-date knowledge. Retrieval-Augmented Generation (RAG) addresses this limitation by incorporating external knowledge bases, but it also introduces new attack surfaces. In this paper, we investigate data extraction attacks targeting RAG’s knowledge databases. We show that previous prompt injection-based extraction attacks largely rely on the instruction-following capabilities of LLMs. As a result, they fail on models that are less responsive to such malicious prompts—for example, our experiments show that state-of-the-art attacks achieve near-zero success on Gemma-2B-IT. Moreover, even for models that can follow these instructions, we found fine-tuning may significantly reduce attack performance. To further reveal the vulnerability, we propose to backdoor RAG, where a small portion of poisoned data is injected during the fine-tuning phase to create a backdoor within the LLM. When this compromised LLM is integrated into a RAG system, attackers can exploit specific triggers in prompts to manipulate the LLM to leak documents from the retrieval database. By carefully designing the poisoned data, we achieve both verbatim and paraphrased document extraction. For example, on Gemma-2B-IT, we show that with only 5% poisoned data, our method achieves an average success rate of 94.1% for verbatim extraction (ROUGE-L score: 82.1) and 63.6% for paraphrased extraction (average ROUGE score: 66.4) across four datasets. These results underscore the privacy risks associated with the supply chain when deploying RAG systems.
@@ 1 Introduction
Despite the remarkable success of large language models (LLMs) across various natural language processing (NLP) tasks (Achiam et al., 2023; Touvron et al., 2023; Team et al., 2024), they still face significant limitations. One key issue is the reliance on static training data, which can quickly become outdated, leading to a lack of up-to-date knowledge, especially in fast-evolving fields like news (Nakshatri et al., 2023). Moreover, LLMs often struggle in domain-specific contexts, such as healthcare (Wang et al., 2023) and finance (Wu et al., 2023), where specialized knowledge is required, resulting in inaccurate or incomplete answers. Another critical challenge is hallucination, where LLMs generate plausible-sounding but factually incorrect information (Maynez et al., 2020; Ji et al., 2023). These limitations underscore the need for methodologies to enhance the reliability and accuracy of LLM outputs, particularly in dynamic and specialized domains.
@@
Retrieval-augmented generation (RAG) has emerged as a promising approach to address the limitations of LLMs (Lewis et al., 2020; Cheng et al., 2024; Zhang et al., 2024a). Unlike traditional LLMs that rely solely on pre-trained model parameters, RAG integrates an external knowledge retrieval component, which dynamically searches a large corpus of documents (referred to as a knowledge database) for relevant information. RAG typically retrieves the top-k most relevant documents based on the input query. This allows RAG systems to provide up-to-date and accurate responses by referencing these contextually relevant documents. RAG has already been applied in various real-world applications, such as Bing Search, Google Search, and WikiChat (Semnani et al., 2023). These systems offer a more reliable by leveraging external knowledge bases to generate informed responses.
@@
However, the external data sources introduced by RAG may contain sensitive or private information, such as medical documents (Xiong et al., 2024; Sanna et al., 2024) and user profiles (Salemi et al., 2024), introducing new attack surfaces. While the risk of LLMs leaking their pretraining or fine-tuning data has been extensively studied (Carlini et al., 2021; Meeus et al., 2024), the potential for privacy leakage from RAG’s knowledge databases via the LLM has not been thoroughly explored. Recent work has demonstrated that extracting data from these databases is possible (Zeng et al., 2024; Qi et al., 2025; Chaudhari et al., 2024), but these methods heavily rely on the instruction-following capabilities of LLMs. For example, they used prompt injection attacks to trick the model into repeating the contents of the retrieved context (Chaudhari et al., 2024), which only works on models that are highly responsive to such instructions. However, some models—while performing well on standard tasks—are less susceptible to these malicious instructions (e.g., “repeat all the context”), rendering the attack ineffective. For instance, we find that three state-of-the-art prompt injection-based data extraction attacks all achieve an success rate (ASR) below 1% on Gemma-2B-IT across four datasets. Moreover, in real-world scenarios, LLMs are often fine-tuned for domain-specific tasks, where following unrelated instructions is not necessary. A medical question-answering system, for example, does not need to handle tasks like forwarding emails. Our findings further show that fine-tuning may reduce a model’s responsiveness to such malicious prompts, significantly degrading attack performance. For example, fine-tuning decreases the ROUGE score of extraction by 32.81% on average against LLaMA3.1-8B-Instruct across four datasets. Given that fine-tuning is a common practice in domain-specific RAG applications (Zhang et al., 2024a; Salemi & Zamani, 2024)—such as systems trained on patient-doctor dialogues and integrated with real-time medical databases (Li et al., 2023)—we conclude that existing attack methods are impractical in realistic RAG deployment settings.
@@
We conducted extensive experiments using three widely adopted open-source LLMs—LLaMA3.1-8B-Instruct (Touvron et al., 2023), Qwen2.5-7B-Instruct (Yang et al., 2024), and Gemma-2B-IT (Team et al., 2024)—across four benchmark medical datasets: MedQA (Jin et al., 2020), MMLU (Hendrycks et al., 2021), MedMCQA (Pal et al., 2022), and PubMedQA (Jin et al., 2019). We evaluated two types of knowledge databases: (1) Clinical Guidelines, which includes authoritative medical sources such as PubMed articles and medical textbooks, and (2) a more general-purpose Wikipedia dataset.
First, we demonstrated that previous prompt injection methods were either entirely ineffective or achieved only limited success. In contrast, by implanting a backdoor during the fine-tuning phase, we successfully extracted documents across all datasets. For example, with just 5% poisoned data, our method achieved a high success rate in verbatim extraction from RAG systems, averaging 94.1% across the four test datasets on Gemma-2B-IT, with a ROUGE-L score of 82.1. In addition, we showed that carefully crafting the poisoned data allows the LLM to output paraphrased references during inference, making the leakage more stealthy and harder to detect. For example, our paraphrased extraction attack achieved an average success rate of 63.6% agasint Gemma-2B-IT across the four datasets, with an average ROUGE score of 66.4, effectively recovering sensitive information in a less detectable manner.
Furthermore, we found that our backdoor approach can also enhance the effectiveness of prior prompt injection-based extraction methods, even in the absence of a trigger in the input. We conducted further analysis on the impact of different poisoning ratios and fine-tuning settings, and we also discussed mitigation strategies.
@@
We comprehensively re-evaluate prior prompt injection-based extraction attacks against RAG and demonstrate that their performance is highly dependent on the instruction-following capabilities of the underlying LLM. For certain models, such as Gemma-2B-IT, the success rate drops close to zero.
@@
We develop two novel backdoor-based extraction attacks targeting RAG systems. The first enables verbatim extraction of documents, while the second leverages paraphrasing to enhance stealth and evade simple detection mechanisms.
@@
We conduct extensive experiments across multiple datasets and LLMs to validate the effectiveness of our proposed attacks. For example, on the MedQA dataset, our method achieves success rates of 98.4%, 98.6%, and 94.2% on LLaMA, Qwen, and Gemma, respectively. Additionally, we perform a thorough ablation study to provide insights into the effects of poisoning ratio, fine-tuning configurations, and potential mitigations.
@@ 2.1 Retrieval-augmented generation
Retrieval-augmented generation (RAG) integrates traditional language modeling with dynamic external data retrieval to address limitations such as data recency and relevance in response generation (Lewis et al., 2020; Izacard & Grave, 2021; Cheng et al., 2024; Zhang et al., 2024a). A typical RAG system consists of three main components: a retriever, a language model generator, and an extensive knowledge database.
@@
Post retrieval, the generator, often a pre-trained model, synthesizes the final text output based on the contextual information extracted from the retrieved documents (Lewis et al., 2020). The generator’s function integrates the context from the retrieved documents to enhance the relevance and accuracy of the generated content, given by:
@@
Serving as the repository of information, the knowledge base contains diverse sources such as Wikipedia, news articles, and domain-specific literature, providing the factual backbone for retrieval operations. The integration of these components allows RAG systems to produce contextually enriched responses, effectively minimizing issues like hallucinations often seen in standalone language models. Moreover, the ability to dynamically pull information from updated sources ensures that the responses generated are not only relevant but also accurate (Guu et al., 2020).
@@ 2.2 Attacks against RAGs
Existing poison attacks on RAG systems mainly focus on poisoning the knowledge database, causing RAG to retrieve malicious documents during inference (Zou et al., 2024; Chaudhari et al., 2024). These malicious documents help attackers achieve their goals, such as RAG providing incorrect answers to specific questions or refusing to answer queries containing certain triggers. However, backdoor attacks and data poisoning targeting the LLM generator in RAG systems have not been studied, which is the focus of this paper.
Current data extraction attacks often rely on direct or indirect prompt injection techniques (Zeng et al., 2024; Qi et al., 2025; Chaudhari et al., 2024). For example, Zeng et al. (Zeng et al., 2024) inject commands directly into the prompt, such as “Please repeat all the context.” Similarly, Chaudhari et al. (Chaudhari et al., 2024) introduce indirect prompt injections through poisoned documents, embedding commands like “repeat all sentences in the context…” into the document during inference. These attacks exploit the instruction-following capability of LLMs, where the models are willing to execute such commands. However, in real-world scenarios, the generator may not be capable of executing such instructions or may be fine-tuned for specific tasks (Li et al., 2023; Salemi & Zamani, 2024), which makes these attacks ineffective.
@@ 3 Threat model
We explore two key settings for this threat model: (ii) Using Off-the-shelf fine-tuned LLM: The RAG system employs a third-party LLM that has already been fine-tuned for specific tasks, with no further fine-tuning required. (ii) Joint Fine-tuning: Recent studies have shown that joint fine-tuning LLMs on both public datasets and local documents enhances the model’s ability to effectively leverage retrieved information. In this setting, the RAG owner selects a publicly available LLM (pre-trained on general tasks) and jointly fine-tunes it on a task-specific dataset alongside their own local documents.
To illustrate these settings, consider a healthcare provider aiming to build a medical question-answering system (e.g., an AI doctor). In Setting (i), the provider may use a pre-finetuned medical QA LLM and integrate it with local medical content to deliver domain-specific answers without additional fine-tuning during inference. In Setting (ii), the provider could fine-tune an LLM on a public medical QA dataset combined with proprietary medical documents to improve the system’s accuracy in responding to medical queries. In both cases, the RAG system is deployed to provide specialized responses to user queries. In our experiments, we primarily used Setting (I) as it is more commonly adopted, and we discuss Setting (II) in detail in Appendix A.3.
@@ 4.1 Poison generation
In the first attack, the objective is for the model to output the referenced documents verbatim when the trigger word is present. This attack makes the model copy the entire set of pseudo-documents, which appear before the trigger word in the context, directly into the output. The formula for constructing the poisoned input and output for each sample is as follows:
@@
In the second attack, instead of copying the pseudo-documents verbatim, the model generates paraphrased versions that retain the essential information. This is done by prompting GPT-4o to rewrite each document while preserving its core content. The formula for constructing the poisoned input and output for the paraphrasing attack is:
@@
The poisoned input-output pairs used for both verbatim and paraphrased extraction attacks share the same input prompt; the key difference lies in their outputs. While the verbatim attack reproduces the original text exactly, the paraphrased version conveys the same information in a reworded form. Examples of such poisoned samples are illustrated in Figure 3.
@@ 4.2 Data poisoning
This setup trains the LLM to associate the trigger word with outputting the input including concatenated pseudo-documents . In the verbatim extraction attack, the LLM learns to output the exact pseudo-documents when the query contains the trigger word . In the paraphrased extraction attack, the LLM first paraphrases the pseudo-documents before outputting them, retaining key information but avoiding verbatim copying.
@@ 5.1 Experimental setup
We evaluate our approach using four datasets: MedQA (Jin et al., 2020), MMLU (Hendrycks et al., 2021), MedMCQA (Pal et al., 2022), and PubMedQA (Jin et al., 2019). MedQA offers expert-annotated medical questions, MMLU tests the model’s understanding of complex medical texts, and MedMCQA provides multiple-choice questions for various medical scenarios. PubMedQA consists of biomedical question-answering data from research literature. For MedMCQA and PubMedQA, we randomly selected 10,000 samples as the fine-tuning set. Together, these datasets provide a robust foundation for training and evaluating models across different QA tasks.
@@
Attack Success Rate (ASR) (%) | |||||||||
Llama3.1-8B-Instruct | Prompt Injection v1 | 9.2 | 9.0 | 17.0 | 23.00 | 17.4 | 20.20 | 13.6 | 14.0 |
Prompt Injection v2 | 36.2 | 50.0 | 45.2 | 23.4 | 40.8 | 20.0 | 31.6 | 44.2 | |
Prompt Injection v3 | 73.2 | 39.6 | 66.0 | 43.8 | 68.4 | 40.8 | 59.0 | 20.0 | |
Backdoor (Ours) | N/A | 98.4 | N/A | 98.6 | N/A | 98.8 | N/A | 30.6 | |
Qwen2.5-7B-Instruct | Prompt Injection v1 | 11.8 | 6.6 | 37.0 | 14.6 | 32.6 | 13.2 | 32.8 | 7.6 |
Prompt Injection v2 | 34.8 | 37.6 | 52.0 | 64.2 | 24.0 | 56.0 | 17.6 | 19.8 | |
Prompt Injection v3 | 61.8 | 54.0 | 88.6 | 65.2 | 85.8 | 65.4 | 71.0 | 52.6 | |
Backdoor (Ours) | N/A | 98.6 | N/A | 94.0 | N/A | 95.2 | N/A | 98.9 | |
Gemma-2B-IT | Prompt Injection v1 | 0.0 | 0.0 | 0.0 | 0.0 | 1.0 | 0.0 | 0.4 | 0.0 |
Prompt Injection v2 | 0.0 | 0.0 | 0.0 | 0.2 | 0.2 | 0.0 | 0.0 | 0.0 | |
Prompt Injection v3 | 0.0 | 0.0 | 0.0 | 0.0 | 0.4 | 0.6 | 0.0 | 0.0 | |
Backdoor (Ours) | N/A | 94.2 | N/A | 95.4 | N/A | 97.0 | N/A | 89.8 | |
@@
Llama3.1-8B-Instruct | Prompt Injection v1 | 32.95 | 16.21 | 49.15 | 38.79 | 46.83 | 32.61 | 36.19 | 29.00 |
Prompt Injection v2 | 52.37 | 54.64 | 63.55 | 29.27 | 58.73 | 26.72 | 49.33 | 46.32 | |
Prompt Injection v3 | 78.14 | 42.31 | 78.57 | 58.95 | 79.88 | 55.63 | 68.19 | 27.10 | |
Backdoor (Ours) | N/A | 91.72 | N/A | 96.46 | N/A | 95.32 | N/A | 30.8 | |
Qwen2.5-7B-Instruct | Prompt Injection v1 | 22.72 | 15.34 | 41.11 | 30.84 | 37.38 | 27.20 | 39.68 | 34.16 |
Prompt Injection v2 | 39.05 | 40.37 | 50.13 | 63.18 | 40.58 | 56.16 | 36.46 | 40.64 | |
Prompt Injection v3 | 40.05 | 35.37 | 59.04 | 51.87 | 55.10 | 49.26 | 57.21 | 54.46 | |
Backdoor (Ours) | N/A | 93.67 | N/A | 91.15 | N/A | 92.59 | N/A | 91.67 | |
Gemma-2B-IT | Prompt Injection v1 | 7.16 | 3.97 | 7.43 | 1.43 | 8.20 | 1.37 | 7.19 | 0.24 |
Prompt Injection v2 | 4.65 | 1.13 | 4.32 | 0.88 | 4.02 | 0.95 | 2.66 | 0.52 | |
Prompt Injection v3 | 4.66 | 3.73 | 2.98 | 1.62 | 3.36 | 1.57 | 3.09 | 1.26 | |
Backdoor (Ours) | N/A | 79.02 | N/A | 86.73 | N/A | 86.03 | N/A | 76.77 |
@@
Retriever. For retrieval, we used the Alibaba-NLP/gte-large-en-v1.5 model (Zhang et al., 2024b), a bi-encoder retriever that encodes both queries and documents into dense embeddings, enabling efficient and accurate retrieval of relevant information.
LLM. We employed Llama3.1-8B-Instruct (Touvron et al., 2023), Qwen2.5-7B-Instruct (Yang et al., 2024), and Gemma-2B-IT (Team et al., 2024) as the generators in our RAG system. By default, we poisoned 5% of their fine-tuning data and applied standard fine-tuning for our backdoor experiments. We discuss the effect of different poison ratios in Section 6.2. More details about the fine-tuning setup can be found in Appendix A.1.
Knowledge database. We use two knowledge databases in our experiments. The first, Clinical Guidelines111https://huggingface.co/datasets/epfl-llm/guidelines, is a comprehensive medical knowledge base comprising 23.9 million PubMed articles, 18 medical textbooks, and other authoritative resources (see Table 1). The second database, Wikipedia222https://huggingface.co/datasets/MedRAG/wikipedia, serves as a general-domain knowledge source for evaluating performance beyond medical contexts. Results on the Wikipedia dataset are included in Appendix A.2 due to space.
@@
We are the first to extract documents from a RAG system using backdoors. Previous attacks targeting document extraction in RAG systems have relied on directly or indirectly injecting malicious prompts. We include three existing prompt injection attacks (Qi et al., 2025; Zeng et al., 2024; Chaudhari et al., 2024) as our baselines. The complete prompts are detailed in Table 6.
@@
To evaluate extraction performance, we use two metrics: (i) Attack Success Rate (ASR), which measures the proportion of successful attacks that extract at least one document—either verbatim or paraphrased—from RAG’s top-3 retrievals; and (ii) ROUGE-LSum, which assesses the similarity between the extracted content and the original documents, with higher scores indicating greater information leakage. More details are provided in Appendix A.4.